The nurse has an order to administer Dopamine at 2 mg/kg/min for a patient admitted to the cardiac intensive care unit with cardiogenic shock. The pharmacy prepares a solution of 400mg/250mL 5% Dextrose. The patient weighs 132 lbs.

The nurse should set the pump at what rate in mL per hour?
Answer #1

Given

Order:Dopamine

Weight:132lbs(0.454×132=59.9 or 60kg)

Dose:2mg/kg/min

As 1kg=2mg

Then 60kg=??mg

Cross multiply and divide

=60×2÷1

=180mg/min

Available:400mg/250ml 5% Dextrose

So to get the total volume

When 400mg=250ml

Then 180mg=??ml

Cross multiply and divide

=180×250÷400

=45000÷400

=112.5 or 113 ml

=113ml

Flow rate=Total volume (ml)÷Hour(hr)

=113÷1

113ml/hr
Answer #2
2mcg/kg/min X 60min/hr X 132lbs X 1kg/2.2lbs X 1mcg/1000mg X 250ml/400mg =3,960,000ml/880,000hr =4.5ml/hr
answered by: Chinenye Onyeka